Robb Fry, Department of Mathematics and Computer Science, Saint Francis Xavier University Antigonish, NS B2G 1C0, Canada, E-mail address:

SMOOTH APPROXIMATION ON BANACH SPACES

We consider the general problem of uniformly approximating continuous functions on Banach spaces by functions with a higher degree of smoothness.

We shall discuss some of the key classical results, as well as more recent and less known developments.